### v1

@@ Abstract
Context:Global first-principles models of pulsar magnetospheres have shown weak particle acceleration in the bulk of the polar caps for low obliquity rotators. This conclusion would lead to the inability of such rotators to generate coherent radio emission, which does not agree with observations. Charge starvation introduced by general-relativistic (GR) effects may explain the pulsar radio emission along the magnetic axis.
Aims:Ab initio global particle-in-cell (PIC) simulations of compact neutron star magnetospheres in the align rotator configuration to investigate the role of GR and plasma supply on the polar cap particle acceleration efficiency - precursor of coherent radio emission.
Methods:A new module for the PIC code OSIRIS to model plasma dynamics around compact objects with full self-consistent GR effects is presented. A detailed description of the extensions and implementation methods are provided for the main sub-algorithms of the PIC loop, including the field solver, particle pusher and charge-conserving current-deposit scheme.
Results:Leptons are efficiently accelerated in the polar caps of neutron stars with force-free magnetospheres. This solution supports strong poloidal currents, which are easily turned to spacelike at any stellar compactness due to the GR frame-dragging effect. Charge-separated magnetospheric solutions, in opposition, depend on the plasma supply and compactness to activate the polar cap. The furthest away the solution is from force-free, the highest the compactness is required for the polar cap activation.
Conclusions:GR effects are crucial to explain the pulsar mechanism for low obliquity rotators. Focusing on the aligned rotator, we show that GR relaxes the minimum required poloidal magnetospheric current for the transition of the polar cap to the accelerator regime, thus justifying the observation of weak pulsars beyond the expected death line. Also, we demonstrate how the interplay between the polar cap and outer gaps might explain the intermittent behaviour of the measured spin-down luminosity and the existence of radio sub-pulse nullings for older pulsars.
OSIRIS-GR: General relativistic activation of the polar cap of a compact neutron star
R. Torres 11    T. Grismayer 11    F. Cruz 11    R.A. Fonseca 1122    L.O. Silva 11 rui.t.torres@tecnico.ulisboa.pt
(Received – – – —- / Accepted – – —-)
Key Words.:
gravitation – magnetic fields – plasmas – methods: numerical – pulsars: general – stars: neutron
offprints: R. Torres,
@@ 1 Introduction
Neutron stars are exquisite physics laboratories coupling extreme magnetic fields with strong gravitational fields. These compact astrophysical objects were discovered in 1967 as radio pulsars (Hewish et al. 1968). Since then have been observed as bright sources of electromagnetic radiation, from radio to gamma rays. Despite recent advances in understanding the intricate dynamics of their magnetospheres, propelled in part by the available computational power, we still lack a complete picture that explains the observed radiation features.
@@
Goldreich & Julian (1969) noted that an isolated neutron star modelled by a rotating spheric perfect conductor that is threaded by a dipolar magnetic field would induce an electric field strong enough to extract charged particles from its surface, populating its magnetosphere with a dense charge-separated plasma. In an ultra-strong magnetic field, the electromagnetic force dominates, leading to the force-free condition, where plasma shorts out the electric field parallel to the magnetic field. Consequently, particles move along magnetic field lines, which corotate rigidly with the neutron star. A crucial element in the magnetosphere is the light cylinder radius, $R_{\mathrm{LC}}=c/\Omega_{*}$, the distance from the rotational axis after which corotation is not possible due to its tangential speed exceeding the light speed. This boundary separates the closed corotating field lines from the open bundle originating from the polar cap of the neutron star. The outflowing particles generate poloidal currents that sweep back the open field lines, inducing a toroidal component. Beyond the light cylinder radius, an equatorial current sheet forms to support the discontinuity of the toroidal magnetic field component. In this region, the poloidal electric field and the toroidal magnetic field result in an outflowing Poynting flux that we observe as a spin-down luminosity. Contopoulos et al. (1999) confirmed this model by numerically solving the ”pulsar equation” (Scharlemann & Wagoner 1973; Michel 1973), the time-independent version of force-free electrodynamics. This work established the standard magnetospheric picture, providing the complete magnetospheric current closure with a y-shaped current layer flowing between the open and closed field lines.
@@
Observations of pulsar wind nebulae (PWNe), i.e. rotation-powered structures formed from the ejection of plasma advected from the inner pulsar magnetosphere (Kirk et al. 2009; Mitchell & Gelfand 2022), support the force-free solution given that there is a dense plasma to screen the non-ideal electric field everywhere (Gruzinov 2005; Spitkovsky 2006; Kalapotharakos, C. & Contopoulos, I. 2009; Timokhin 2007; Bai & Spitkovsky 2010; Pétri 2012; Philippov & Spitkovsky 2014). However, these models do not constrain the mechanism that provides the plasma to drive the required magnetospheric currents. Without external plasma sources, there must exist dissipative regions - gaps - within the magnetosphere where particles accelerate and produce pairs that cancel the accelerating electric field. These gaps would operate at much shorter time scales (microseconds) than the magnetospheric time scales (milliseconds to seconds, thus quasi-stationary), consistent with the force-free solution. Several studies explored the observational implications of the polar cap gap (Daugherty & Harding 1982; Usov & Melrose 1995; Daugherty & Harding 1996; Ruderman & Sutherland 1975), slot gap (Arons & Scharlemann 1979; Arons 1983; Muslimov & Harding 2003, 2004), and outer gap models (Cheng et al. 1986; Romani 1996; Takata et al. 2004; Hirotani 2008), highlighting the importance of these gaps not only to provide the necessary plasma to the magnetosphere but also to explain the observed electromagnetic spectrum.
@@
The radio beam in radio pulsars originates from a region above the pair formation front of the polar cap gap due to the nonlinear screening of the accelerating electric field. The misalignment of the magnetic and rotation axis leads to a pulsed behaviour as the beam crosses our line of sight periodically, similar to a lighthouse. These radio pulses contain microstructures, some possessing single (Drake & Craft 1968; Deshpande & Rankin 1999; Vivekanand & Joshi 1999) or multiple drifting components (Qiao et al. 2004; Basu et al. 2018). Models to explain this behaviour assume a limitation in the extracted charged particles from the stellar surface such that the spark filaments no longer corotate with it. Hence, the sub-pulses undergo an $E\times B$ drift different from the field lines due to the parallel electric field at the base of those field lines. Ruderman & Sutherland (1975) (hereafter RS75) assumed the case where no particles could be lifted from the stellar surface, thus forming a vacuum gap above it. The intermediate model of Gil, J. et al. (2003) allows a thermionic outflow from the surface forming a partially screened gap (PSG). This model accommodates a necessary reduction of the drifting speed and provides a better estimate of the X-ray luminosity from the backflow bombardment of particles. The space charge limited flow model (SCLF) by Arons & Scharlemann (1979), where particles can freely escape from the stellar surface, does not provide a natural explanation for this phenomenon. The binding energies of either ions or electrons should not be too high, thus motivating the SCLF model (Medin & Lai 2010). Alternatively, studies have highlighted the possibility of strongly bound particles in bare strange quark stars (Ren-xin & Guo-jun 1998; Xu et al. 1999; Yu & Xu 2011), favouring the RS75 or PSG models.
@@
First principle global magnetospheric models that capture the kinetic plasma scales are the best way to introduce the effects of these dissipative regions. Several works have proposed tools to tackle this very computationally demanding task in one-dimension (Timokhin 2010; Timokhin & Arons 2012), two-dimensions (Chen & Beloborodov 2014; Belyaev 2015b, a; Cerutti et al. 2013, 2015; Philippov et al. 2015b; Cerutti, B. & Philippov, A. A. 2017; Cruz et al. 2021; Hu & Beloborodov 2022; Bransgrove et al. 2022; Cruz et al. 2023) or three-dimensions (Philippov et al. 2015a; Cerutti & Beloborodov 2016; Philippov & Spitkovsky 2018; Kalapotharakos et al. 2018; Brambilla et al. 2018; Cerutti, Benoît et al. 2020; Hakobyan et al. 2023). The work by Philippov et al. (2015a) concluded that the polar cap of low obliquity rotators has weak particle acceleration, suppressing the pair production and shutting down the radio pulsar mechanism. Several authors later introduced the general-relativistic frame-dragging effect (Beskin 1990; Muslimov & Tsygan 1992) into Maxwell’s equations, demonstrating that this effect could potentiate the ignition of the polar cap even for the aligned rotator (Belyaev & Parfrey 2016; Gralla et al. 2016; Philippov et al. 2015b; Huang et al. 2018; Chen et al. 2020; Philippov et al. 2020; Bransgrove et al. 2022). However, ab initio simulations of neutron star magnetospheres with full self-consistent general-relativistic effects still lack in the community.
@@
The purpose of this paper is to present simulations of compact neutron stars including self-consistent general-relativistic effects. Also, we compactify in a single publication all the necessary techniques required for the implementation of a GR-PIC code, which does not exist systematically in the literature, as far as we know. We detail the methodology and algorithms used to extend the particle-in-cell code OSIRIS to include GR effects in Sect. 2. In Sects. 3 and 4, we perform numerical experiments of such magnetospheres by varying the plasma injection methods to explore the full range of magnetospheric solutions, from quasi-force-free to full-charge-separated solutions, expanding on the consequent observational implications. Conclusions and future prospects are presented in Sect. 5.
@@ 2 GR-PIC method
In this section, we detail the extension module of general relativity to the OSIRIS particle-in-cell framework (Fonseca et al. 2002). We describe the main equations and integration algorithms implemented in OSIRIS including the field solver, equations of motion, charge conservation, and boundary conditions. A general description of the techniques and methodologies (independent of the internal specifics/implementation of the PIC code) is provided to facilitate their implementation in other PIC codes.
Throughout the paper, we use units in which $c=G=1$, the (,+,+,+)(-,+,+,+)( - , + , + , + ) metric signature, and stellar properties are characterized by an asterisk in subscript (e.g., $R_{*}$ for the stellar radius). In addition, we adopt the 3+1 formalism of general relativity (Thorne & MacDonald 1982).
@@ 2.1 Metric
The astrophysical objects we are interested in constitute the class of compact objects where the background metric is no longer a flat (Minkowski) metric. We focus our study on the dynamics of the magnetospheres of black holes and neutron stars. The electromagnetic energy density in such environments is negligibly small compared to their corresponding total mass density. Consequently, we can assume a fixed background metric for these systems. Black holes and neutron stars are modeled by the Kerr and the Hartle-Thorne metric (Hartle & Thorne 1968), respectively. However, as neutron stars rotate at significantly smaller angular velocities, one can invoke the slow-rotation approximation and simplify the corresponding metric ($R_{*}\Omega_{*}\ll 1$) for neutron stars. In this approximation, and keeping only the linear terms of the angular velocity, the exterior metric takes the form (Hartle 1967):
$\mathrm{d}s^{2}=$ $-\alpha^{2}\mathrm{d}t^{2}+\@tens{\gamma}_{ij}\left(\mathrm{d}x^{i}+\beta^{i}\mathrm{d}t\right)\left(\mathrm{d}x^{j}+\beta^{j}\mathrm{d}t\right)$
@@
where $\alpha=\sqrt{1-R_{s}/r}$ and $\beta^{i}=\left(0,0,-\omega\right)$ are the lapse function and shift vector, respectively, ${\gamma}_{ij}$ is the spatial part of the metric, $R_{s}$ is the Schwarzschild radius, and $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}{\theta}\mathrm{d}\phi^{2}$. Interestingly, this metric is equivalent to the slow-rotation approximation of the Kerr metric, which leads us to adopt the Boyer-Lindquist coordinate system ($t$,$r$,$\theta$,$\phi$). Currently, OSIRIS has the Minkowski, the Schwarzschild, and the slow-rotation approximation of the Kerr metric implemented. The Minkowski metric is recovered in the zero compactness limit ($R_{s}\rightarrow 0$), and the Schwarzschild metric by neglecting the frame-dragging effect ($\omega(r)\rightarrow 0$). The latter effect is a correction that captures the differential rotation $\omega(r)$ associated with a free-falling inertial frame (Ravenhall & Pethick 1994):
@@
We define fiducial observers (FIDOs) with world lines normal to spatial hypersurfaces of constant ”universal” time. The lapse function relates the rate at which the local fiducial observer clock ticks in relation to the universal time. Hence, the lapse function translates local time-measured quantities to the universal time. The shift vector measures the angular velocity at which the coordinate frame shears in relation to the FIDOs. We select FIDOs corotating with the absolute space, thus considering zero angular momentum observers (ZAMOs). In these hypersurfaces, we use the orthonormal basis vectors and vectorial components, i.e. $e_{\hat{i}}\equiv e_{i}/\sqrt{\gamma_{ii}}$ and $A^{\hat{i}}\equiv\sqrt{\gamma_{ii}}A^{i}$ such that ${A}=A^{i}e_{i}=A^{\hat{i}}e_{\hat{i}}$. This basis allows for a trivial extension of three-dimensional vectorial operations to curved geometries. Also, the definition of line, area, and volume elements is modified according to eq. (1) - see Appendix A.

### v2

@@ Abstract
We present ab initio global general-relativistic Particle-in-cell (GR-PIC) simulations of compact millisecond neutron star magnetospheres in the axisymmetric aligned rotator configuration. We investigate the role of GR and plasma supply on the polar cap particle acceleration efficiency - the precursor of coherent radio emission - employing a new module for the PIC code OSIRIS, designed to model plasma dynamics around compact objects with fully self-consistent GR effects. We provide a detailed description of the main sub-algorithms of the novel PIC algorithm, including a charge-conserving current deposit scheme for curvilinear coordinates. We demonstrate efficient particle acceleration in the polar caps of compact neutron stars with denser magnetospheres, numerically validating the spacelike current extension provided by force-free models. We show that GR relaxes the minimum required poloidal magnetospheric current for the transition of the polar cap to the accelerator regime, thus justifying the observation of weak pulsars beyond the expected death line. We denote that spin-down luminosity intermittency and radio pulse nullings for older pulsars might arise from the interplay between the polar and outer gaps. Also, narrower radio beams are expected for weaker low-obliquity pulsars.
keywords:
General relativity , Stellar magnetospheres , Plasma astrophysics , Computational methods , Compact objects , Millisecond pulsars
journal: New Astronomy
[…]
[first]organization=GoLP/Instituto de Plasmas e Fusão Nuclear,addressline=Instituto Superior Técnico, Universidade de Lisboa, postcode=1049-001, state=Lisboa, country=Portugal
[…]
[second]organization=DCTI/ISCTE,addressline=Instituto Universitário de Lisboa, postcode=1649-026, state=Lisboa, country=Portugal
@@ 1 Introduction
Neutron stars are exquisite physics laboratories coupling extreme magnetic fields with strong gravitational fields. These compact astrophysical objects were discovered in 1967 as radio pulsars (Hewish et al., 1968). Since then pulsars have been observed as bright sources of electromagnetic radiation, from radio to gamma rays. Despite recent advances in understanding the intricate dynamics of their magnetospheres, propelled in part by the available computational power, we still lack a complete picture that explains the observed radiation features.
@@
Goldreich and Julian (1969) noted that an isolated neutron star modelled by a rotating spheric perfect conductor that is threaded by a dipolar magnetic field would induce an electric field strong enough to extract charged particles from its surface, populating its magnetosphere with a dense charge-separated plasma. In an ultra-strong magnetic field, the electromagnetic force dominates, leading to the force-free condition, where plasma shorts out the electric field parallel to the magnetic field. Consequently, particles move along magnetic field lines, which corotate rigidly with the neutron star. A crucial element in the magnetosphere is the light cylinder radius, $R_{\mathrm{LC}}=c/\Omega_{*}$, the distance from the rotational axis after which corotation is not possible due to its tangential speed exceeding the light speed. This boundary separates the closed corotating field lines from the open bundle originating from the polar cap of the neutron star. The outflowing particles generate poloidal currents that sweep back the open field lines, inducing a toroidal component. Beyond the light cylinder radius, an equatorial current sheet forms to support the discontinuity of the toroidal magnetic field component. In this region, the poloidal electric field and the toroidal magnetic field result in an outflowing Poynting flux that we observe as a spin-down luminosity. Contopoulos et al. (1999) confirmed this model by numerically solving the “pulsar equation” (Scharlemann and Wagoner, 1973; Michel, 1973), the time-independent version of force-free electrodynamics. This work established the standard magnetospheric picture, providing the complete magnetospheric current closure with a y-shaped current layer flowing between the open and closed field lines.
@@
Observations of pulsar wind nebulae (PWNe), i.e. rotation-powered structures formed from the ejection of plasma advected from the inner pulsar magnetosphere (Kirk et al., 2009; Mitchell and Gelfand, 2022), support the force-free solution given that there is a dense plasma to screen the non-ideal electric field everywhere (Gruzinov, 2005; Spitkovsky, 2006; Kalapotharakos, C. and Contopoulos, I., 2009; Timokhin, 2007; Bai and Spitkovsky, 2010; Pétri, 2012; Philippov and Spitkovsky, 2014). However, these models do not constrain the mechanism that provides the plasma to drive the required magnetospheric currents. Without external plasma sources, there must exist dissipative regions - gaps - within the magnetosphere where particles accelerate and produce pairs that cancel the accelerating electric field. These gaps would operate at much shorter time scales (microseconds) than the magnetospheric time scales (milliseconds to seconds, thus quasi-stationary), consistent with the force-free solution. Several studies explored the observational implications of the polar cap gap (Daugherty and Harding, 1982; Usov and Melrose, 1995; Daugherty and Harding, 1996; Ruderman and Sutherland, 1975), slot gap (Arons and Scharlemann, 1979; Arons, 1983; Muslimov and Harding, 2003, 2004), and outer gap models (Cheng et al., 1986; Romani, 1996; Takata et al., 2004; Hirotani, 2008), highlighting the importance of these gaps not only to provide the necessary plasma to the magnetosphere but also to explain the observed electromagnetic spectrum (e.g., Jun-Tao et al. (2021)).
@@
The radio beam in radio pulsars originates from a region above the pair formation front of the polar cap gap due to the nonlinear screening of the accelerating electric field. The misalignment of the magnetic and rotation axis leads to a pulsed behaviour as the beam crosses our line of sight periodically, similar to a lighthouse. These radio pulses contain microstructures, some possessing single (Drake and Craft, 1968; Deshpande and Rankin, 1999; Vivekanand and Joshi, 1999) or multiple drifting components (Qiao et al., 2004; Basu et al., 2018). Models to explain this behaviour assume a limitation in the extracted charged particles from the stellar surface such that the spark filaments no longer corotate with it. Hence, the sub-pulses undergo an $E\times B$ drift different from the field lines due to the parallel electric field at the base of those field lines. Ruderman and Sutherland (1975) (hereafter RS75) assumed the case where no particles could be lifted from the stellar surface, thus forming a vacuum gap above it. The intermediate model of Gil, J. et al. (2003) allows a thermionic outflow from the surface forming a partially screened gap (PSG). This model accommodates a necessary reduction of the drifting speed and provides a better estimate of the X-ray luminosity from the backflow bombardment of particles. The space charge limited flow model (SCLF) by Arons and Scharlemann (1979), where particles can freely escape from the stellar surface, does not provide a natural explanation for this phenomenon. The binding energies of either ions or electrons should not be too high, thus motivating the SCLF model (Medin and Lai, 2010). Alternatively, studies have highlighted the possibility of strongly bound particles in bare strange quark stars (Ren-xin and Guo-jun, 1998; Xu et al., 1999; Yu and Xu, 2011), favouring the RS75 or PSG models.
@@
First principle global magnetospheric models that capture the kinetic plasma scales are the best way to introduce the effects of these dissipative regions. Several works have proposed tools to tackle this very computationally demanding task in one-dimension (Timokhin, 2010; Timokhin and Arons, 2012), two-dimensions (Chen and Beloborodov, 2014; Belyaev, 2015b, a; Cerutti et al., 2013, 2015; Philippov et al., 2015a; Cerutti, B. and Philippov, A. A., 2017; Cruz et al., 2021; Hu and Beloborodov, 2022; Bransgrove et al., 2022; Cruz et al., 2023) or three-dimensions (Philippov et al., 2015b; Cerutti and Beloborodov, 2016; Philippov and Spitkovsky, 2018; Kalapotharakos et al., 2018; Brambilla et al., 2018; Cerutti, Benoît et al., 2020; Hakobyan et al., 2023). The work by Philippov et al. (2015b) concluded that the polar cap of low obliquity rotators has weak particle acceleration, suppressing the pair production and shutting down the radio pulsar mechanism. Several authors later introduced the general-relativistic frame-dragging effect (Beskin, 1990; Muslimov and Tsygan, 1992) into Faraday’s induction equation, demonstrating that this effect could potentiate the ignition of the polar cap even for the aligned rotator (Philippov et al., 2015a; Chen et al., 2020; Philippov et al., 2020; Bransgrove et al., 2022) as predicted by general-relativistic force-free models (Belyaev and Parfrey, 2016; Gralla et al., 2016; Huang et al., 2018). Although such kinetic models are able to capture the polar cap ignition, qualitative corrections from neglecting the lapse function and general-relativistic magnetic field topology (Torres et al., 2023) are not resolved. One of such parameters of utmost importance is the spacelike angle $\theta_{\mathrm{SL}}$, expected to limit the poloidal extension of the observed pulsar radio emission generation volume. This angle was obtained analytically for force-free models (Belyaev and Parfrey, 2016; Gralla et al., 2016), but was not yet verified using general-relativistic particle-in-cell simulations. Thus, ab initio simulations of neutron star magnetospheres with fully self-consistent general-relativistic effects still lack in the community.
@@
The purpose of this paper is to present the first simulations of compact millisecond neutron stars including self-consistent general-relativistic effects, using a charge-conserving GR-PIC code. Also, we summarize all the necessary techniques required for the implementation of a GR-PIC code. As far as we know, a systematic description of the GR-PIC algorithm does not exist in the literature. We detail the methodology and algorithms used to extend the particle-in-cell code OSIRIS to include GR effects in Sect. 2. In Sects. 3 and 4, we perform numerical experiments of such magnetospheres by varying the plasma injection methods to explore the full range of magnetospheric solutions, from quasi-force-free to full-charge-separated solutions, expanding on the consequent observational implications. We show that the correct dimensions of the polar cap and spacelike current region can only be obtained when considering the lapse function effects. Also, we demonstrate that charge-separated magnetospheric solutions possess smaller values of the spacelike angle with respect to force-free (FF) estimates. Therefore, we conjecture the existence of narrower radio beams from weak pulsars, some still observable beyond the expected death line. Conclusions and future prospects are presented in Sect. 5.
@@ 2 GR-PIC method
In this section, we detail the extension module of general relativity to the OSIRIS particle-in-cell framework (Fonseca et al., 2002). We describe the main equations and integration algorithms implemented in OSIRIS including the field solver, equations of motion, charge conservation, and boundary conditions. A general description of the techniques and methodologies (independent of the internal specifics/implementation of the PIC code) is provided to facilitate their implementation in other PIC codes.
Throughout the paper, we use units in which $c=G=1$, the (,+,+,+)(-,+,+,+)( - , + , + , + ) metric signature, and stellar properties are characterized by an asterisk in subscript (e.g., $R_{*}$ for the stellar radius). In addition, we adopt the 3+1 formalism of general relativity (Thorne and MacDonald, 1982).
@@ 2.1 Metric
The astrophysical objects we are interested in constitute the class of compact objects where the background metric is no longer a flat (Minkowski) metric. We focus our study on the dynamics of the magnetospheres of black holes and neutron stars. The electromagnetic energy density in such environments is negligibly small compared to their corresponding total mass density. Consequently, we can assume a fixed background metric for these systems. Black holes and neutron stars are modeled by the Kerr and the Hartle-Thorne metric (Hartle and Thorne, 1968), respectively. However, as neutron stars rotate at significantly smaller angular velocities, one can invoke the slow-rotation approximation and simplify the corresponding metric ($R_{*}\Omega_{*}\ll 1$) for neutron stars. In this approximation, and keeping only the linear terms of the angular velocity, the exterior metric takes the form (Hartle, 1967):
$\mathrm{d}s^{2}=$ $-\alpha^{2}\mathrm{d}t^{2}+{\gamma}_{ij}\left(\mathrm{d}x^{i}+\beta^{i}\mathrm{d}t\right)\left(\mathrm{d}x^{j}+\beta^{j}\mathrm{d}t\right)$
@@
where $\alpha=\sqrt{1-R_{s}/r}$ and $\beta^{i}=\left(0,0,-\omega\right)$ are the lapse function and shift vector, respectively, ${\gamma}_{ij}$ is the spatial part of the metric, $R_{s}$ is the Schwarzschild radius, and $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}{\theta}\mathrm{d}\phi^{2}$. Interestingly, this metric is equivalent to the slow-rotation approximation of the Kerr metric, which leads us to adopt the Boyer-Lindquist coordinate system ($t$,$r$,$\theta$,$\phi$). Currently, OSIRIS has the Minkowski, the Schwarzschild, and the slow-rotation approximation of the Kerr metric implemented. The Minkowski metric is recovered in the zero compactness limit ($R_{s}\rightarrow 0$), and the Schwarzschild metric by neglecting the frame-dragging effect ($\omega(r)\rightarrow 0$). The latter effect is a correction that captures the differential rotation $\omega(r)$ associated with a free-falling inertial frame (Ravenhall and Pethick, 1994):
@@
We define fiducial observers (FIDOs) with world lines normal to spatial hypersurfaces of constant “universal” time. The lapse function relates the rate at which the local fiducial observer clock ticks in relation to the universal time. Hence, the lapse function translates local time-measured quantities to the universal time. The shift vector measures the angular velocity at which the coordinate frame shears in relation to the FIDOs. We select FIDOs corotating with the absolute space, thus considering zero angular momentum observers (ZAMOs). In these hypersurfaces, we use the orthonormal basis vectors and vectorial components, i.e. $e_{\hat{i}}\equiv e_{i}/\sqrt{\gamma_{ii}}$ and $A^{\hat{i}}\equiv\sqrt{\gamma_{ii}}A^{i}$ such that $\boldsymbol{A}=A^{i}e_{i}=A^{\hat{i}}e_{\hat{i}}$. This basis allows for a trivial extension of three-dimensional vectorial operations to curved geometries. Also, the definition of line, area, and volume elements is modified according to eq. (1) - see A.